Thomas Private Limited Company is about to start production of new products to

sell to customers. Before the start of the new production, it costs the company $

200,000 to construct the factory’s building. Once the construction is over, the cost

per unit of production is estimated to be birr 20. Upon selling, the company will

incur a 20% commission expense. Moreover, the price per unit of products is

decided to be Birr 50.

Required:

a)

Construct the total cost equation in terms of quantity.

b)

Determine the break-even quantity (Revenue)


1
Expert's answer
2022-03-27T11:04:40-0400

a) "TC=TFC+VC*Q=200,000*0.2+20Q=40+20Q"

b)"BEQ={FC \over P-VC}={40 \over 50-20}=1.3333"